Determine IB,IC,and VCE for a base biased transistor circuit with the following values: ẞ= 90, Vcc = 12V, RB= 22kΩ and RC= 100Ω
1
Expert's answer
2020-10-02T07:13:21-0400
  • For input circuit, we have the equation-

IB = "\frac{V_{CC}-0.7}{R_B}"

Here, VCC is also base biasing voltage.

"\frac{I_C}{\beta}" = "\frac{12-0.7}{22000}"

Given "\beta" = 90

IC = "\frac{101.7}{2200}"

IC = 46.23 mA ...(1)

  • For output circuit, we have the equation-

12 = ICRC + VCE

12 = (0.04623)(100) + VCE

VCE = 7.38 V

  • Also, IB = "\frac{46.23}{90}"

IB = 0.514 mA